Liquid bromine (Br2) has a density of 2.91g/mL. How many molecules of bromine are contained in a 8.50-mL sample of bromine?
1
Expert's answer
2020-08-06T03:55:30-0400

m = V×ρ

m(Br2) = 8.50 mL × 2.91 g/mL = 24.735 g

M(Br2) = 159.808 g/mol

n = m/M

n(Br2) = 24.735 g / 159.808 (g/mol) = 0.154 mol

One mole of Br2 molecules contains 6.02×1023 Br2 molecules.

Number of molecules = n×6.02×1023 = 0.154×6.02×1023 = 9.27×1022 molecules.

Answer: 9.27×1022 molecules